A chemical engineer must calculate the maximum safe operating temperature of a high-pressure gas reaction vessel. The vessel is a stainless-steel cylinder that measures 20.0cm wide and 24.0cm high. The maximum safe pressure inside the vessel has been measured to be 4.50MPa. For a certain reaction the vessel may contain up to .235kg of dinitrogen monoxide gas. Calculate the maximum safe operating temperature the engineer should recommend for this reaction. Write your answer in degrees Celsius. Be sure your answer has the correct number of significant digits.
Answer #1

Given : Pressure of a N2O gas = 4.50 MPa

Width of cylinder = 20.0 cm

Height of cylinder = 24.0 cm

Mass of a N2O gas = 0.235 kg = 235 g

Maximum safe operating Temperature (T) = ?

We have relation, P V = n R T

where, P is a pressure of a gas, V is a volume of a gas, n is no. of moles of gas, R is a gas constant and T is temperature of gas.

We need to calculate Volume & number of moles of N2O.

In case of gases , volume of gas is equal to volume container. Volume of container is calculated by using formula r 2 h , where r is radius & h is height of cylinder.

Radius of cylinder = Width / 2 = 20.0 cm / 2 = 10.0 cm

Volume of N2O gas = volume of cylinder = ( 10.0 cm ) 2 24.0 cm = 7539.8 cm 3

We have relation, 1 L = 1000 cm 3

Volume of N2O gas = 7539.8 cm 3 ( 1 L / 1000 cm 3 ) =7.5398 L

We know that, No. of moles = Mass / Molar mass

No. of moles of N2O gas = 235 g / 44.01 g/ mol = 5.340 mol

Pressure of N2O gas = 4.50 MPa = 4.50 10 6 Pa

We have relation, 1 atm = 101325 Pa

Pressure of N2O gas =4.50 10 6 Pa ( 1 atm / 101325 Pa ) = 44.41 atm

We have relation, P V = n R T

T = P V / n R

Maximum safe operating Temperature (T) = 44.41 atm 7.5398 L / 5.340 mol 0.082057 L atm mol -1 K -1  

Maximum safe operating Temperature (T) = 764.16 K

Maximum safe operating Temperature (T) =764.16 - 273 = 491.16 0 C = 491 0 C

ANSWER : Maximum safe operating Temperature (T) the engineer should recommend for this reaction is 491 0 C.